# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 6. Understanding nestedness through connectivity
The purpose of this section is to characterise the property of total-nestedness for tetra-separations $(A,B)$ in terms of connectivity-properties of their separators $S(A,B)$; see 6.16. This translates the problem of proving total-nestedness of $(A,B)$ to a problem about constructing various families of paths that link up the vertices and edges in the separator $S(A,B)$ in different ways.
For an element $x\in V(G)\cup E(G)$ we let $\hat{x}$ denote the singleton $\{x\}$ if $x$ is a vertex, and the set of endvertices of $x$ if $x$ is an edge. For a set $X\subseteq V(G)\cup E(G)$ we write ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\hat{X}}}:=\bigcup_{x\in X}\hat{x}$.
@@
Let $G$ be a $k$-connected graph. Then $G$ has no $-\ell$-separation for any $\ell<k$.
@@
Assume for a contradiction that $(A,B)$ is a $-\ell$-separation of $G$ with $\ell<k$. Then $S(A,B)$ contains an edge $ab$ with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. We claim that $A\smallsetminus(B\cup\{a\})$ or $B\smallsetminus(A\cup\{b\})$ is non-empty. Indeed, otherwise $V(G)=(A\cap B)\cup\{a,b\}$, so $|G|=\ell+1\leqslant k$, contradicting that $k$-connected graphs such as $G$ have $>k$ vertices by definition. So suppose that there is a vertex $v\in A\smallsetminus(B\cup\{a\})$. Let $X$ be obtained from $S(A,B)$ by keeping all vertices, replacing $ab$ with $a$, and replacing each other edge with its endvertex in $B\smallsetminus A$. Then $X$ is an $\ell$-separator of $G$ that separates $v$ from $b$, contradicting that $G$ is $k$-connected. ∎
@@
    the maximum number of independent $X$$Y$ paths in $G$;
@@
As usual, (i) $\leqslant$ (ii), so it remains to show (i) $\geqslant$ (ii). Since every $X$$Y$ edge forms a path in every maximum set of independent $X$$Y$ paths in $G$, we may assume without loss of generality that $G$ contains no $X$$Y$ edges. Let $X^{\prime}:=N_{G}(X)$ and $Y^{\prime}:=N_{G}(Y)$. Note that $X^{\prime}\cup Y^{\prime}$ is disjoint from $X\cup Y$. Let $G^{\prime}:=G-(X\cup Y)$.
[…]
The ‘moreover’-part follows from Lemma 6.1. ∎
@@ 6.1. Half-connected
A mixed-separation $(A,B)$ of a graph $G$ is half-connected if $G[A\smallsetminus B]$ or $G[B\smallsetminus A]$ is connected [Tridecomp].
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $(A,B)$ is not half-connected; that is, $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ are not connected. Then $S(A,B)$ consists of vertices only. Let $C_{A}$ be a component of $G[A\smallsetminus B]$ and let $C_{B}$ be a component of $G[B\smallsetminus A]$. Let $C:=V(C_{A})\cup V(C_{B})\cup S(A,B)$ and $D:=V(G-C_{A}-C_{B})$. Then $(C,D)$ is a genuine 4-separation with $S(C,D)=S(A,B)$ that crosses $(A,B)$ with all links empty. To see that $(C,D)$ is a tetra-separation, note first that both $G[C\smallsetminus D]$ and $G[D\smallsetminus C]$ are disconnected. Since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in every component of $G-S(A,B)$. So $(C,D)$ satisfies the degree-condition, and it trivially satisfies the matching-condition. ∎
@@
Let $(C^{\prime\prime},D^{\prime\prime})$ be the left-right-shift of $(C^{\prime},D^{\prime})$, which is a tetra-separation by 5.10. First, note that $\hat{x}_{1}\subseteq C^{\prime}\smallsetminus D^{\prime}\subseteq C^{\prime\prime}\smallsetminus D^{\prime\prime}$ and $\hat{x}_{2}\subseteq D^{\prime}\smallsetminus C^{\prime}\subseteq D^{\prime\prime}\smallsetminus C^{\prime\prime}$ by 5.9. So, $x_{1}$ and $x_{2}$ are in the $C^{\prime\prime}$-link and the $D^{\prime\prime}$-link, respectively. In particular, $(C^{\prime\prime},D^{\prime\prime})$ crosses $(A,B)$.
@@ 6.3. Potter-linked
Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a graph $G$. Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see Figure 4.2. Then $\{u_{2},u_{3}u_{4}\}$ is $A$-weird and $\{u_{3},u_{1}u_{2}\}$ is $C$-weird.
Let $X,Y\subseteq V(G)\cup E(G)$. An $X$$Y$ edge means an edge $xy$ with $x\in X$ and $y\in Y$.
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 0-potter-linked. Then there are at most four independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$. By (Menger for independent paths). (6.2), there is a mixed-4-separation $(C,D)$ of $G$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\subseteq C\smallsetminus D$ and $\hat{\pi}_{2}\subseteq D\smallsetminus C$. In particular, $|C\smallsetminus D|\geqslant 2$ and $|D\smallsetminus C|\geqslant 2$. Observe that $(C,D)$ crosses $(A,B)$ since the $C$-link and the $D$-link are non-empty. Let $(C^{\prime},D^{\prime})$ be the left-right-shift of $(C,D)$, which is a tetra-separation by 5.10.
@@
The following definition is supported by Figure 6.1.
@@
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links, say $\pi_{1}$ is the $C$-link and $\pi_{2}$ is the $D$-link, and so that exactly one corner is potter, say the $AC$-corner. We show that $\pi_{1}$ is not 1-potter-linked to $\pi_{2}$ with respect to $(A,B)$. For this, we show that for $(X,Y)=(A,B)$ in Definition 6.9, all of (6.9.1)(6.9.3) fail.
Observe that $\pi_{1}$ is $A$-weird by Example 6.6, so (6.9.1) fails. By assumption, the corner $AC$ is the only corner that is potter. By the (Crossing Lemma). (4.1), every link has size two while the centre is empty. Moreover, there do not exist jumping or diagonal edges, and the $B$-link contains no dangling edge. It follows that there is no $\pi_{1}$$(\hat{\pi}_{2}\cap B)$ edge and no $(\hat{\pi}_{1}\cap B)$$\pi_{2}$ edge, so (6.9.2) fails. Moreover, every $\hat{\pi}_{1}$$\hat{\pi}_{2}$ path in $G[B]$ has an internal vertex or an edge in the $B$-link. Thus, there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$, so (6.9.3) fails.
@@
By $\neg$ (6.9.3), there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$. By (Menger for independent paths). (6.2), there is a mixed-separation $(C_{1},D_{1})$ of $G[B]$ of order $\leqslant 2$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\cap B\subseteq C_{1}\smallsetminus D_{1}$ and $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}$. In particular, $|C_{1}\smallsetminus D_{1}|\geqslant 2$ and $|D_{1}\smallsetminus C_{1}|\geqslant 2$. See Figure 6.2 for the next definition. We define $(C_{2},D_{2}):=(C_{1}\cup a_{1},(D_{1}\cup A)\smallsetminus\{v_{1}\})$. We have $S(C_{2},D_{2})=S(C_{1},D_{1})\cup\{a^{\prime}_{1}v_{1},a_{1}\}$ since $v_{1}$ does not have a neighbour in $A$ besides $a_{1}^{\prime}$. By 4-connectivity of $G$ and Lemma 6.1, $(C_{2},D_{2})$ has order exactly 4. Let $(C,D)$ be the left-right-shift of $(C_{2},D_{2})$. Since $C_{1}\smallsetminus D_{1}\subseteq C_{2}\smallsetminus D_{2}$ we have $|C_{2}\smallsetminus D_{2}|\geqslant 2$, and similarly $|D_{2}\smallsetminus C_{2}|\geqslant 2$. Hence $(C,D)$ is a tetra-separation by 5.10.
@@
Consider the crossing-diagram for $(A,B)$ and $(C,D)$. We claim that the $C$-link includes $\pi_{1}$. The $C$-link clearly contains $v_{1}$. The vertex $a_{1}\in S(C_{2},D_{2})$ has two neighbours $v_{1},b_{1}$ in $C_{2}\smallsetminus D_{2}$, so $a_{1}\in S(C,D)$. Hence $e_{1}\notin S(C,D)$. Since the end $b_{1}$ of $e_{1}$ lies in $C_{1}\smallsetminus D_{1}\subseteq C\smallsetminus D$ by 5.9, the edge $e_{1}$ lies in the $C$-link. Next, we claim that the $D$-link includes $\pi_{2}$. Every vertex in $\pi_{2}$ clearly lies in the $D$-link. Now let $f$ be an edge in $\pi_{2}$. The endvertex of $f$ in $B\smallsetminus A$ lies in $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}\subseteq D\smallsetminus C$ by 5.9. The endvertex of $f$ in $A\smallsetminus B$ must be distinct from $a_{1}$ by the matching-condition, and hence is contained in $D_{2}\smallsetminus C_{2}\subseteq D\smallsetminus C$ by 5.9. Therefore, both ends of $f$ lie in $D\smallsetminus C$ while $f\in S(A,B)$, so $f$ lies in the $D$-link.
@@
Since the $C$-link includes $\pi_{1}$ and the $D$-link includes $\pi_{2}$, the two tetra-separations $(A,B)$ and $(C,D)$ cross, and the inclusions are equalities by the (Crossing Lemma). (4.1). It remains to show that the $AC$-corner is the only corner that is potter. The edge $v_{1}a_{1}^{\prime}$ dangles from the $C$-link through the $A$-link, so the corner $AC$ is potter by the (Crossing Lemma). (4.1). By 4.7, the corners for $AD$ and $BC$ are not potter. Assume for contradiction that the $BD$-corner is potter. Then $\pi_{2}$ is $B$-weird and hence contains an edge $f$. But then the endvertex of $f$ in $B\smallsetminus A$ is contained in $D\smallsetminus C$ as shown above. In particular, the corner for $BD$ is not empty, contradicting that $BD$ is potter. ∎
Figure 6.2. The situation in the second half of the proof of Lemma 6.10. $\pi_{1}$ is red and $\pi_{2}$ is blue.
@@
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links, say $\pi_{1}$ is the $C$-link and $\pi_{2}$ is the $D$-link, and so that exactly two corners are potter. By 4.7, no two adjacent corners are potter. So we can assume without loss of generality that the $AC$-corner and the $BD$-corner are potter while the other two corners are not potter. Then $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird by Example 6.6. So, (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 2-potter-linked with respect to $(A,B)$. Say (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$; that is to say, $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird. For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see Figure 6.3. Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A\smallsetminus\{a_{1}\}$, and let $b_{2}^{\prime}$ be the unique neighbour of $v_{2}$ in $B\smallsetminus\{b_{1}\}$. By the degree-condition, we have $a^{\prime}_{1}\in A\smallsetminus B$ and $b^{\prime}_{1}\in B\smallsetminus A$.
Let $C:=B\smallsetminus\{v_{2}\})\cup\{a_{1}\}$ and $D:=(A\smallsetminus\{v_{1}\})\cup\{b_{2}\}$. Since $\pi_{1}$ is $A$-weird, $\pi_{2}$ is $B$-weird, and $v_{1}v_{2}\notin E(G)$ by 6.12, we have $S(C,D)=\{a^{\prime}_{1}v_{1},a_{1},b_{2},v_{2}b^{\prime}_{2}\}$. Note that $(C,D)$ crosses $(A,B)$ since $v_{1}$ is in the $C$-link and $v_{2}$ is in the $D$-link. The vertex $a_{1}$ has two neighbours in $C\smallsetminus D$, namely $v_{1}$ and $b_{1}$. Since $a_{1}$ has degree at least four and it is not adjacent to $b_{2}$, the only other vertex in $S(C,D)$, it follows that $a_{1}$ also has at least two neighbours in $D\smallsetminus C$. Similarly, $b_{2}$ also has two neighbours in both $C\smallsetminus D$ and $D\smallsetminus C$. Hence $(C,D)$ satisfies the degree-condition. It satisfies the matching-condition since $v_{1}\neq v_{2}$ means that $a^{\prime}_{1}v_{1}$ and $v_{2}b^{\prime}_{2}$ do not share endvertices. Hence, $(C,D)$ is a tetra-separation that crosses $(A,B)$ with $\pi_{1}$ and $\pi_{2}$ as opposite links. The corners $AC$ and $BD$ are potter by construction. By 4.7, no other corner is potter. ∎
Figure 6.3. The situation in the second half of the proof of Lemma 6.13. $\pi_{1}$ is red and $\pi_{2}$ is blue.
@@
By 4.7, every crossing pair of tetra-separations has at most two corners that are potter. So the corollary immediately follows from Lemma 6.8, Lemma 6.10 and Lemma 6.13. ∎
@@
We combine Lemma 6.3, Lemma 6.5 and Corollary 6.14 with the (Crossing Lemma). (4.1). ∎

### v2

@@ 6. Understanding nestedness through connectivity
The purpose of this section is to characterise the property of total-nestedness for tetra-separations $(A,B)$ in terms of connectivity-properties of their separators $S(A,B)$; see ˜6.16. This translates the problem of proving total-nestedness of $(A,B)$ to a problem about constructing various families of paths that link up the vertices and edges in the separator $S(A,B)$ in different ways.
For an element $x\in V(G)\cup E(G)$ we let $\hat{x}$ denote the singleton $\{x\}$ if $x$ is a vertex, and the set of endvertices of $x$ if $x$ is an edge. For a set $X\subseteq V(G)\cup E(G)$ we write $^{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\hat{X}}}:=\bigcup_{x\in X}\hat{x}$.
@@
Let $G$ be a $k$-connected graph. Then $G$ has no mixed-$\ell$-separation for any $\ell<k$.
@@
Assume for a contradiction that $(A,B)$ is a mixed-$\ell$-separation of $G$ with $\ell<k$. Then $S(A,B)$ contains an edge $ab$ with $a\in A\smallsetminus B$ and $b\in B\smallsetminus A$. We claim that $A\smallsetminus(B\cup\{a\})$ or $B\smallsetminus(A\cup\{b\})$ is non-empty. Indeed, otherwise $V(G)=(A\cap B)\cup\{a,b\}$, so $|G|=\ell+1\leqslant k$, contradicting that $k$-connected graphs such as $G$ have >k>k vertices by definition. So suppose that there is a vertex $v\in A\smallsetminus(B\cup\{a\})$. Let $X$ be obtained from $S(A,B)$ by keeping all vertices, replacing $ab$ with $a$, and replacing each other edge with its endvertex in $B\smallsetminus A$. Then $X$ is an $\ell$-separator of $G$ that separates $v$ from $b$, contradicting that $G$ is $k$-connected. ∎
@@
    the maximum number of independent XXYY paths in $G$;
@@
As usual, (i) $\leqslant$ (ii), so it remains to show (i) $\geqslant$ (ii). Since every XXYY edge forms a path in every maximum set of independent XXYY paths in $G$, we may assume without loss of generality that $G$ contains no XXYY edges. Let $X^{\prime}:=N_{G}(X)$ and $Y^{\prime}:=N_{G}(Y)$. Note that $X^{\prime}\cup Y^{\prime}$ is disjoint from $X\cup Y$. Let $G^{\prime}:=G-(X\cup Y)$.
Now let $\mathcal{P}$ be a maximum-size set of independent XXYY paths in $G$. Then $\mathcal{P}$ determines a set $\mathcal{P}^{\prime}$ of $|\mathcal{P}|$ pairwise disjoint $X^{\prime}$$Y^{\prime}$ paths in $G^{\prime}$. Hence $G^{\prime}$ contains an $X^{\prime}$$Y^{\prime}$ separator $S^{\prime}$ of size $|\mathcal{P}|$, by Menger’s theorem applied in $G^{\prime}$ to $X^{\prime},Y^{\prime}$ and $\mathcal{P}^{\prime}$. Then $S^{\prime}$ is the separator of a mixed-separation of $G$ that strongly separates $X$ and $Y$, and which has order $|S^{\prime}|=|\mathcal{P}|$.
The ‘moreover’-part follows from Lemma˜6.1. ∎
@@ 6.1. Half-connected
A mixed-separation $(A,B)$ of a graph $G$ is half-connected if $G[A\smallsetminus B]$ or $G[B\smallsetminus A]$ is connected [CK23].
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $(A,B)$ is not half-connected; that is, $G[A\smallsetminus B]$ and $G[B\smallsetminus A]$ are not connected. Then $S(A,B)$ consists of vertices only. Let $C_{A}$ be a component of $G[A\smallsetminus B]$ and let $C_{B}$ be a component of $G[B\smallsetminus A]$. Let $C:=V(C_{A})\cup V(C_{B})\cup S(A,B)$ and $D:=V(G-C_{A}-C_{B})$. Then $(C,D)$ is a genuine 4-separation with $S(C,D)=S(A,B)$ that crosses $(A,B)$ with all links empty. To see that $(C,D)$ is a tetra-separation, note first that both $G[C\smallsetminus D]$ and $G[D\smallsetminus C]$ are disconnected. Since $G$ is 4-connected, every vertex in $S(A,B)$ has a neighbour in every component of GS(A,B)G-S(A,B). So $(C,D)$ satisfies the degree-condition, and it trivially satisfies the matching-condition. ∎
@@
Let $(C^{\prime\prime},D^{\prime\prime})$ be the left-right-reduction of $(C^{\prime},D^{\prime})$, which is a tetra-separation by ˜5.10. First, note that $\hat{x}_{1}\subseteq C^{\prime}\smallsetminus D^{\prime}\subseteq C^{\prime\prime}\smallsetminus D^{\prime\prime}$ and $\hat{x}_{2}\subseteq D^{\prime}\smallsetminus C^{\prime}\subseteq D^{\prime\prime}\smallsetminus C^{\prime\prime}$ by ˜5.9. So, $x_{1}$ and $x_{2}$ are in the $C^{\prime\prime}$-link and the $D^{\prime\prime}$-link, respectively. In particular, $(C^{\prime\prime},D^{\prime\prime})$ crosses $(A,B)$.
@@ 6.3. Potter-linked
Let $(A,B)$ and $(C,D)$ be crossing tetra-separations in a graph $G$. Assume that the corner for $AC$ is potter, witnessed by an edge $u_{1}u_{2}$ that dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and by an edge $u_{3}u_{4}$ that dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link; see Figure˜4.2. Then $\{u_{2},u_{3}u_{4}\}$ is $A$-weird and $\{u_{3},u_{1}u_{2}\}$ is $C$-weird.
Let $X,Y\subseteq V(G)\cup E(G)$. An XXYY edge means an edge $xy$ with $x\in X$ and $y\in Y$.
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 0-potter-linked. Then there are at most four independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$. By (Menger for independent paths). (6.2), there is a mixed-4-separation $(C,D)$ of $G$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\subseteq C\smallsetminus D$ and $\hat{\pi}_{2}\subseteq D\smallsetminus C$. In particular, $|C\smallsetminus D|\geqslant 2$ and $|D\smallsetminus C|\geqslant 2$. Observe that $(C,D)$ crosses $(A,B)$ since the $C$-link and the $D$-link are non-empty. Let $(C^{\prime},D^{\prime})$ be the left-right-reduction of $(C,D)$, which is a tetra-separation by ˜5.10.
@@
The following definition is supported by Figure˜6.1.
@@
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links, say $\pi_{1}$ is the $C$-link and $\pi_{2}$ is the $D$-link, and so that exactly one corner is potter, say the $AC$-corner. We show that $\pi_{1}$ is not 1-potter-linked to $\pi_{2}$ with respect to $(A,B)$. For this, we show that for $(X,Y)=(A,B)$ in Definition˜6.9, all of (6.9.1)(6.9.3) fail.
Observe that $\pi_{1}$ is $A$-weird by Example˜6.6, so (6.9.1) fails. By assumption, the corner $AC$ is the only corner that is potter. By the (Crossing Lemma). (4.1), every link has size two while the centre is empty. Moreover, there do not exist jumping or diagonal edges, and the $B$-link contains no dangling edge. It follows that there is no $\pi_{1}$$(\hat{\pi}_{2}\cap B)$ edge and no $(\hat{\pi}_{1}\cap B)$$\pi_{2}$ edge, so (6.9.2) fails. Moreover, every $\hat{\pi}_{1}$$\hat{\pi}_{2}$ path in $G[B]$ has an internal vertex or an edge in the $B$-link. Thus, there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$, so (6.9.3) fails.
@@
By $\neg$ (6.9.3), there are at most two independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[B]$. By (Menger for independent paths). (6.2), there is a mixed-separation $(C_{1},D_{1})$ of $G[B]$ of order $\leqslant 2$ that strongly separates $\hat{\pi}_{1}$ from $\hat{\pi}_{2}$, say with $\hat{\pi}_{1}\cap B\subseteq C_{1}\smallsetminus D_{1}$ and $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}$. In particular, $|C_{1}\smallsetminus D_{1}|\geqslant 2$ and $|D_{1}\smallsetminus C_{1}|\geqslant 2$. See Figure˜6.2 for the next definition. We define $(C_{2},D_{2}):=(C_{1}\cup a_{1},(D_{1}\cup A)\smallsetminus\{v_{1}\})$. We have $S(C_{2},D_{2})=S(C_{1},D_{1})\cup\{a^{\prime}_{1}v_{1},a_{1}\}$ since $v_{1}$ does not have a neighbour in $A$ besides $a_{1}^{\prime}$. By 4-connectivity of $G$ and Lemma˜6.1, $(C_{2},D_{2})$ has order exactly 4. Let $(C,D)$ be the left-right-reduction of $(C_{2},D_{2})$. Since $C_{1}\smallsetminus D_{1}\subseteq C_{2}\smallsetminus D_{2}$ we have $|C_{2}\smallsetminus D_{2}|\geqslant 2$, and similarly $|D_{2}\smallsetminus C_{2}|\geqslant 2$. Hence $(C,D)$ is a tetra-separation by ˜5.10.
@@
Consider the crossing-diagram for $(A,B)$ and $(C,D)$. We claim that the $C$-link includes $\pi_{1}$. The $C$-link clearly contains $v_{1}$. The vertex $a_{1}\in S(C_{2},D_{2})$ has two neighbours $v_{1},b_{1}$ in $C_{2}\smallsetminus D_{2}$, so $a_{1}\in S(C,D)$. Hence $e_{1}\notin S(C,D)$. Since the end $b_{1}$ of $e_{1}$ lies in $C_{1}\smallsetminus D_{1}\subseteq C\smallsetminus D$ by ˜5.9, the edge $e_{1}$ lies in the $C$-link. Next, we claim that the $D$-link includes $\pi_{2}$. Every vertex in $\pi_{2}$ clearly lies in the $D$-link. Now let $f$ be an edge in $\pi_{2}$. The endvertex of $f$ in $B\smallsetminus A$ lies in $\hat{\pi}_{2}\cap B\subseteq D_{1}\smallsetminus C_{1}\subseteq D\smallsetminus C$ by ˜5.9. The endvertex of $f$ in $A\smallsetminus B$ must be distinct from $a_{1}$ by the matching-condition, and hence is contained in $D_{2}\smallsetminus C_{2}\subseteq D\smallsetminus C$ by ˜5.9. Therefore, both ends of $f$ lie in $D\smallsetminus C$ while $f\in S(A,B)$, so $f$ lies in the $D$-link.
@@
Since the $C$-link includes $\pi_{1}$ and the $D$-link includes $\pi_{2}$, the two tetra-separations $(A,B)$ and $(C,D)$ cross, and the inclusions are equalities by the (Crossing Lemma). (4.1). It remains to show that the $AC$-corner is the only corner that is potter. The edge $v_{1}a_{1}^{\prime}$ dangles from the $C$-link through the $A$-link, so the corner $AC$ is potter by the (Crossing Lemma). (4.1). By ˜4.7, the corners for $AD$ and $BC$ are not potter. Assume for contradiction that the $BD$-corner is potter. Then $\pi_{2}$ is $B$-weird and hence contains an edge $f$. But then the endvertex of $f$ in $B\smallsetminus A$ is contained in $D\smallsetminus C$ as shown above. In particular, the corner for $BD$ is not empty, contradicting that $BD$ is potter. ∎
Figure 6.2. The situation in the second half of the proof of Lemma˜6.10. $\pi_{1}$ is red and $\pi_{2}$ is blue.
@@
$\neg$ (ii) $\Rightarrow$ $\neg$ (i). Assume that a tetra-separation $(C,D)$ crosses $(A,B)$ so that $\pi_{1}$ and $\pi_{2}$ form opposite links, say $\pi_{1}$ is the $C$-link and $\pi_{2}$ is the $D$-link, and so that exactly two corners are potter. By ˜4.7, no two adjacent corners are potter. So we can assume without loss of generality that the $AC$-corner and the $BD$-corner are potter while the other two corners are not potter. Then $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird by Example˜6.6. So, (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $\{\pi_{1},\pi_{2}\}$ is not 2-potter-linked with respect to $(A,B)$. Say (6.11.1) and (6.11.2) fail for $(X,Y)=(A,B)$; that is to say, $\pi_{1}$ is $A$-weird and $\pi_{2}$ is $B$-weird. For $i\in\{1,2\}$, write $\pi_{i}=:\{v_{i},e_{i}\}$ where $v_{i}$ is a vertex and $e_{i}=a_{i}b_{i}$ is an edge with $a_{i}\in A\smallsetminus B$ and $b_{i}\in B\smallsetminus A$; see Figure˜6.3. Let $a^{\prime}_{1}$ be the unique neighbour of $v_{1}$ in $A\smallsetminus\{a_{1}\}$, and let $b_{2}^{\prime}$ be the unique neighbour of $v_{2}$ in $B\smallsetminus\{b_{1}\}$. By the degree-condition, we have $a^{\prime}_{1}\in A\smallsetminus B$ and $b^{\prime}_{1}\in B\smallsetminus A$.
Let $C:=B\smallsetminus\{v_{2}\})\cup\{a_{1}\}$ and $D:=(A\smallsetminus\{v_{1}\})\cup\{b_{2}\}$. Since $\pi_{1}$ is $A$-weird, $\pi_{2}$ is $B$-weird, and $v_{1}v_{2}\notin E(G)$ by ˜6.12, we have $S(C,D)=\{a^{\prime}_{1}v_{1},a_{1},b_{2},v_{2}b^{\prime}_{2}\}$. Note that $(C,D)$ crosses $(A,B)$ since $v_{1}$ is in the $C$-link and $v_{2}$ is in the $D$-link. The vertex $a_{1}$ has two neighbours in $C\smallsetminus D$, namely $v_{1}$ and $b_{1}$. Since $a_{1}$ has degree at least four and it is not adjacent to $b_{2}$, the only other vertex in $S(C,D)$, it follows that $a_{1}$ also has at least two neighbours in $D\smallsetminus C$. Similarly, $b_{2}$ also has two neighbours in both $C\smallsetminus D$ and $D\smallsetminus C$. Hence $(C,D)$ satisfies the degree-condition. It satisfies the matching-condition since $v_{1}\neq v_{2}$ means that $a^{\prime}_{1}v_{1}$ and $v_{2}b^{\prime}_{2}$ do not share endvertices. Hence, $(C,D)$ is a tetra-separation that crosses $(A,B)$ with $\pi_{1}$ and $\pi_{2}$ as opposite links. The corners $AC$ and $BD$ are potter by construction. By ˜4.7, no other corner is potter. ∎
Figure 6.3. The situation in the second half of the proof of Lemma˜6.13. $\pi_{1}$ is red and $\pi_{2}$ is blue.
@@
By ˜4.7, every crossing pair of tetra-separations has at most two corners that are potter. So the corollary immediately follows from Lemma˜6.8, Lemma˜6.10 and Lemma˜6.13. ∎
@@
We combine Lemma˜6.3, Lemma˜6.5 and Corollary˜6.14 with the (Crossing Lemma). (4.1). ∎